# Probing the signature of axions through the quasinormal modes of black holes

### v1

@@ Abstract
The advent of gravitational-wave astronomy opened up a new window for probing the physics in strong-gravity regimes [1]. From the merger events of compact binaries, one can constrain not only the masses and charges of black holes (BHs) but also quasinormal modes (QNMs) of damped oscillations. QNMs of the Schwarzschild BH can be modified by the presence of extra degrees of freedom [2, 3, 4, 5, 6, 7], e.g., vector and scalar fields. A simple example is the Reissner-Nordström (RN) BH with an electric charge [8, 9, 10, 11], which arises from the presence of a $U(1)$ gauge-invariant vector field $A_{\mu}$ in Einstein-Maxwell theory.
@@
where $g$ is the determinant of metric tensor $g_{\mu\nu}$, $M_{\rm Pl}$ is the reduced Planck mass, $R$ is the Ricci scalar, and $m_{\phi}$ is the axion mass. The field strength tensor $F_{\mu\nu}$ is related to a $U(1)$ invariant vector field $A_{\mu}$, as $F_{\mu\nu}=\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}$, and $\tilde{F}^{\mu\nu}=\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}/(2\sqrt{-g})$ with $\epsilon^{0123}=+1$.
@@
Since this condition is violated for $q_{M}=0$ or $q_{E}=0$, we need the existence of both magnetic and electric charges to realize a nontrivial axion hair. Without loss of generality, we will consider the case $\phi_{0}>0$, $q_{M}>0$, $q_{E}>0$, and $g_{a\gamma\gamma}<0$. Because of Eq. (8), combining (10) with $h_{1}>0$ gives
@@
In this section, we compute the QNMs of hairy BHs in EMA theory by considering linear perturbations on the background (2). Choosing the Regge-Wheeler-Zerilli gauge [30, 31, 32], the components of metric perturbations $h_{\mu\nu}$ are given by
@@
where $Y_{l}(\theta)$’s are the $m=0$ components of spherical harmonics $Y_{lm}(\theta,\varphi)$, and $Y_{l,\theta}(\theta)\equiv{\rm d}Y_{l}(\theta)/{\rm d}\theta$. In Eq. (16), we omit the summation of $Y_{l}(\theta)$ concerning the multipoles $l$. The vector and axion fields have the following perturbed components
@@
respectively. The coupling $-(\sqrt{-g}/4)g_{a\gamma\gamma}\phi F_{\mu\nu}\tilde{F}^{\mu\nu}$ in the action (1) does not have contributions from the metric components and depends only linearly on each perturbed field. Moreover, none of the perturbations are coupled to the modes with different values of $l$ or $m$. Without loss of generality, we will focus on each multipole mode $l\geq 2$ with $m=0$. Since the perturbations in the odd- and even-parity sectors are mixed for $q_{M}\neq 0$, we must deal with them all at once.
Expanding Eq. (1) up to quadratic order in perturbations, the resulting second-order action contains ten perturbed variables $H_{0}$, $H_{1}$, $H_{2}$, $h_{1}$, $Q$, $W$, $\delta A_{0}$, $\delta A_{1}$, $\delta A$, and $\delta\phi$. The explicit form of the total second-order action ${\cal S}^{(2)}$ is given in Appendix A. Introducing the new fields $\chi_{1}$, $v_{1}$, $\chi_{2}$ defined in Eqs. (27), (28), and (32), respectively, we can express the action in terms of the five dynamical perturbations $\chi_{1}$, $v_{1}$, $\chi_{2}$, $\delta A$, $\delta\phi$, and their $t,r$ derivatives. The process for deriving the reduced second-order action is explained in Appendix B. Here, $v_{1}$ and $\chi_{1}$ are associated with the even- and odd-parity gravitational perturbations, respectively, while $\chi_{2}$ and $\delta A$ arise from the vector-field perturbations in even- and odd-parity sectors, respectively. We also have the dynamical axion perturbation $\delta\phi$.
@@
We consider the two fundamental QNMs that are present in the limit $q_{M}\to 0$. For $q_{M}=0$, the background solution reduces to the electrically charged RN BH without the axion hair. In this case, there are one gravitational and the other electromagnetic QNMs, whose frequencies were computed in Refs. [9, 8, 9]. The study of a possible extra fundamental QNM due to the nontrivial axion profile is left for future work. In fact, in this study, we would like to focus on the crucial property of hairy BHs to distinguish the magnetic charge from the electric one at the level of the gravitational/electromagnetic QNMs, leading to an unequivocal sign for the existence of axions. We choose a configuration with $q_{M}=q_{T}\sin\alpha$ and $q_{E}=q_{T}\cos\alpha$, where the total charge $q_{T}$ is chosen to be $\sqrt{13/50}$. We vary the angle $\alpha$ from 0 to $12/50$, where each solution differs from the previous one by the interval $\Delta\alpha=1/100$. For each value of $\alpha$, we numerically solve the background equations of motion and find the value of $r_{h}$ leading to $M=1$, so that all the BHs have the same mass and total charge.
@@
There are several interesting extensions of our work. First, the computation of QNMs for charged rotating BH solutions with the axion hair [33] is the important next step for placing realistic bounds on our model parameters. Next, the gravitational waveforms emitted during the inspiral phase of charged binary BHs with the axion hair will put further constraints on the theory. Thirdly, the observations of BH shadows such as the Event Horizon Telescope [34] will give upper bounds on the BH charges. Finally, it will be of interest to study the effect of large magnetic fields on the BH physics near the horizon, e.g., restoration of an electroweak symmetry [12]. These issues are left for future work.

### v2

@@ Abstract
The advent of gravitational-wave astronomy opened up a new window for probing the physics in strong-gravity regimes [1]. From the merger events of compact binaries, one can constrain not only the masses and charges of black holes (BHs) but also quasinormal modes (QNMs) of damped oscillations. QNMs of the Schwarzschild BH can be modified by the presence of extra degrees of freedom [2, 3, 4, 5, 6, 7], e.g., vector and scalar fields. A simple example is the Reissner-Nordström (RN) BH with an electric charge [8, 9, 10, 11], which arises from the presence of a vector field $A_{\mu}$ in Einstein-Maxwell theory.
@@
where $g$ is the determinant of metric tensor $g_{\mu\nu}$, $M_{\rm Pl}$ is the reduced Planck mass, $R$ is the Ricci scalar, and $m_{\phi}$ is the axion mass. The field strength tensor $F_{\mu\nu}$ is related to the vector field $A_{\mu}$, as $F_{\mu\nu}=\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}$, and $\tilde{F}^{\mu\nu}=\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}/(2\sqrt{-g})$ with $\epsilon^{0123}=+1$. The action (1) respects $U(1)$ gauge invariance under the shift $A_{\mu}\to A_{\mu}+\nabla_{\mu}\chi$.
@@
Since this condition is violated for $q_{M}=0$ or $q_{E}=0$, we need the existence of both magnetic and electric charges to realize a nontrivial axion hair. The inequality (11) does not hold for $g_{a\gamma\gamma}=0$ either, so we require the axion-photon coupling $-(1/4)g_{a\gamma\gamma}\phi F_{\mu\nu}\tilde{F}^{\mu\nu}$ to realize hairy BH solutions. In other words, the no-hair property of BHs for a canonical scalar field [28, 29] is broken by the appearance of a secondary axion hair through interaction with electromagnetic fields.
Without loss of generality, we will consider the case $\phi_{0}>0$, $q_{M}>0$, $q_{E}>0$, and $g_{a\gamma\gamma}<0$. Because of Eq. (8), combining (10) with $h_{1}>0$ gives
@@
In this section, we compute the QNMs of hairy BHs in EMA theory by considering linear perturbations on the background (2). We choose the gauge in which the $\theta$ and $\varphi$ components of $h_{\mu\nu}$ vanish, i.e.,
@@
where $Y_{l}(\theta)$’s are the $m=0$ components of spherical harmonics $Y_{lm}(\theta,\varphi)$, and $Y_{l,\theta}(\theta)\equiv{\rm d}Y_{l}(\theta)/{\rm d}\theta$. In Eq. (16), we omit the summation of $Y_{l}(\theta)$ concerning the multipoles $l$. Note that the above gauge choice is different from the Regge-Wheeler-Zerilli gauge [32, 33, 34], but the former can also fix the residual gauge degrees of freedom completely [35, 36, 37, 38]. Since our theory has $U(1)$ gauge symmetry, we can choose a gauge in which the $\theta$ component of the vector-field perturbation $\delta A_{\mu}$ vanishes [38]. Then, we consider the perturbed components of vector and axion fields, as
@@
respectively. The coupling $-(\sqrt{-g}/4)g_{a\gamma\gamma}\phi F_{\mu\nu}\tilde{F}^{\mu\nu}$ in the action (1) does not have contributions from the metric components and depends only linearly on each perturbed field. Moreover, none of the perturbations are coupled to the modes with different values of $l$ or $m$. The background spherical symmetry allows us to set $m=0$ without loss of generality. Indeed, for fixed $l$, the second-order perturbed action does not depend on the values of $m$. Since the perturbations in the odd- and even-parity sectors are mixed for $q_{M}\neq 0$, we must deal with them all at once.
For $l\geq 2$, we expand Eq. (1) up to quadratic order in perturbations. Then, the resulting second-order action contains ten perturbed variables $H_{0}$, $H_{1}$, $H_{2}$, $h_{1}$, $Q$, $W$, $\delta A_{0}$, $\delta A_{1}$, $\delta A$, and $\delta\phi$. The explicit form of the total second-order action ${\cal S}^{(2)}$ is given in Appendix A. Introducing the new fields $\chi_{1}$, $v_{1}$, $\chi_{2}$ defined in Eqs. (27), (28), and (32), respectively, we can express the action in terms of the five dynamical perturbations $\chi_{1}$, $v_{1}$, $\chi_{2}$, $\delta A$, $\delta\phi$, and their $t,r$ derivatives. The process for deriving the reduced second-order action is explained in Appendix B. Here, $v_{1}$ and $\chi_{1}$ are associated with the even- and odd-parity gravitational perturbations, respectively, while $\chi_{2}$ and $\delta A$ arise from the vector-field perturbations in even- and odd-parity sectors, respectively. We also have the dynamical axion perturbation $\delta\phi$.
@@
[…]
@@
There are several interesting extensions of our work. First, the computation of QNMs for charged rotating BH solutions with the axion hair [39] is the important next step for placing realistic bounds on our model parameters. Next, the gravitational waveforms emitted during the inspiral phase of charged binary BHs with the axion hair will put further constraints on the theory. Thirdly, the observations of BH shadows such as the Event Horizon Telescope [40] will give upper bounds on the BH charges. Fourth, we leave a detailed study of the isospectrality of QNMs for a future separate work. While isospectrality may be broken, this letter aims to demonstrate the potential for simultaneously finding magnetic charges and axions through the QNMs of BHs. Finally, it will be of interest to study the effect of large magnetic fields on the BH physics near the horizon, e.g., restoration of an electroweak symmetry [12]. These issues are left for future work.